OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1). y = one thirdx − 3 y = one thirdx + 17 y = −3x − 3 y = −3x + 17

OpenStudy (anonymous):

lets find the slope by putting your equation in y = mx + b form, where m is the slope. 3x + y = 7 y = -3x + 7 the slope of this equation is -3 are you with me so far ?

OpenStudy (anonymous):

a perpendicular slope will have a negative reciprocal....that just means you " flip " the fraction over and change the sign......so the negative reciprocal of -3 is 1/3. That will be the slope you need. now we will use the formula: y - y1 = m(x - x1) using slope(m) = 1/3 and points (6,-1) -- x1 = 6 and y1 = -1 now plug our info into the formula y - y1 = m(x - x1) y - (-1) = 1/3(x - 6) can you find the line now ?